OpenStudy (anonymous):

Solve for x in the proportion 9/27x=x/x+3

OpenStudy (anonymous):

Solve for x in the proportion 9/27x = x/x+3 I understand : 9/(27x) = x/(x + 3) 1/(3x) = x/(x + 3) Cross multiply : x + 3 = 3x^2 3x^2 - x - 3 = 0 --------------------------------------… To find the roots for the quadratic equations : Roots = (-b +/- sqrt(b^2 - 4ac)) / 2a Where : a = 3 b = -1 c = -3 Verification of the rationality : sqrt(b^2 - 4ac) sqrt((-1)^2 - (4*3*-3)) sqrt(1 + 36) roots are rationals coz sqrt(b^2-4ac) = 6,083 ---> not negative root 1 = (-b + 6,083) / 2a root 1 = (1 + 6,083) / (2 *3) root 1 = 7,083 / 6 root 1 --------------------> 1,181 root 2 = (-b - 6,083) / 2a root 2 = (1 - 6,083) / (2 *3) root 2 = -5,083 / 6 root 2 --------------------> -0,847 --------------------------------------… So x ≈ 1.18 and x ≈ -0.85 (c)

OpenStudy (anonymous):

set it equal to zero then use the quadratic formula

OpenStudy (anonymous):

woah o.O okay so can you show me a simpler way of doing that. all of that was just alot o.O

OpenStudy (anonymous):

lopez_hatesmath already tok over

OpenStudy (anonymous):

yeahhh okay i gotchu

OpenStudy (anonymous):

9/27x)=x/(x+3) 27^2-9x-27=0 divide by 9 3x^2-x-3=0 Using quadratic formula

OpenStudy (anonymous):

3x^2-x-3=0 plug this in for the quadratic formula formula which is ( x=-b+ 0r -sqrt of b^2-4ac all over (2a)